OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find f(-1) + 5g(-3)

OpenStudy (anonymous):

f(-1)=2(-1)+1 =-2+1=-1 g(-3)=3(-3)+1 =-9+1=-8 -8*5=-40 -1+(-40)=-41